### v1

# Vertex evaluation of multiplex graphs using Forman Curvature

abstract
Identifying vertices that play a central role is a fundamental problem in network analysis. Although traditional centrality measures have been widely used for this purpose, the growing complexity of contemporary networks necessitates more sophisticated indicators. Forman curvature has recently emerged as a promising approach. In this paper, we define Forman curvature for multilayer networks, a class of complex networks characterized by multiple types of connections or layers between nodes, which are increasingly used to model intricate real-world phenomena. We establish the key properties of Forman curvature in the context of multilayer networks and demonstrate its utility for identifying vertices that hold central positions within these networks. Furthermore, we show that Forman curvature can also serve as an effective tool for the structural classification of entire multilayer networks.
1. Introduction
Graph theory provides essential tools for visualizing and analyzing complex systems across the physical and social sciences, as well as information engineering. To enhance realism, graphs are often augmented with weights assigned to vertices and/or edges, resulting in weighted graphs. Extending fundamental concepts such as node degree and distance metrics to weighted graphs introduces complexity, thereby increasing the difficulty of analysis. In particular, analytical methods specifically designed for doubly-weighted graphs, which feature weights on both vertices and edges, remain relatively scarce. However, even doubly-weighted graphs may not fully capture the intricacies of many real-world situations where, depending on the event, multiple weights may be assigned to vertices and edges. For example, consider a trading network modeling relationships between companies. Edges represent transactions, but a comprehensive analysis might require multiple associated weights, such as transaction volume, transportation distance, and carbon emissions. Multilayer graphs offer a framework to address this complexity. The concept and terminology of multilayer networks have been explored by numerous researchers over the years. For instance, Castelblanco et al. [3] employed a multilayer network approach for multidimensional analysis encompassing economic transaction networks and risk propagation networks among stakeholders in megaprojects. The definitions and implementations of these networks vary considerably in the literature. Allard et al. [2] proposed a definition for multilayer or multitype networks based on assigning a type from a set of $M$ types to each node in a set of $N$ nodes. In their comprehensive review, Kivelä et al. [10] cataloged 26 distinct names used for related network structures, noting that while some terms represent the same structure and framework, others denote specific structural variations. Because this study focuses on applications within transportation networks, it adopts the specific formalism of multiplex networks, a subtype of multilayer graphs. Kurant et al. [11] modeled transportation networks and infrastructure using edge-weighted multilevel graphs, demonstrating advantages over conventional models. Similarly, Aleta et al. [1] analyzed the traffic networks of nine European cities using unweighted multilevel graphs, assessing vulnerability and identifying key nodes via centrality computations. Building upon this, we propose a more sophisticated analysis method by incorporating weights on both vertices and edges within the multiplex network framework. Notably, many analysis methods, including that proposed by Aleta et al., cannot be directly applied to such doubly-weighted graphs.
A key challenge in network analysis is finding centrally important vertices. Centrality measures are commonly employed for this task. However, the increasing scale and complexity of network data available today necessitate the development of more advanced analysis methods. Specifically, there is a lack of well-established centrality measures tailored for doubly-weighted graphs, which are the primary focus of this study. The clustering coefficient, another powerful concept in network analysis, also faces limitations. Although Regino et al. [5] introduced and demonstrated the utility of a clustering coefficient for unweighted multilevel graphs, clustering coefficients cannot be defined for graphs featuring weights on both vertices and edges, making them unsuitable for the doubly-weighted multiplex networks considered herein. Therefore, network curvature, derived from differential geometry, has attracted attention as an alternative measure distinct from traditional centrality or clustering metrics.
Classical curvature, defined on smooth Riemannian manifolds to quantify local geometry, is not directly applicable to discrete structures such as graphs. However, discrete curvature, introduced as an extension of the concept of curvature, has proven valuable for characterizing graph structure ([7, 9, 12]). Building on this theoretical foundation, research on applying discrete curvature is being conducted across various domains.
Several distinct definitions of discrete curvature exist, with Ollivier–Ricci curvature (based on optimal transport theory) and Forman–Ricci curvature (based on weight functions) being representative examples. Ollivier-Ricci curvature uses the $L^{1}$-Wasserstein distance from optimal transport theory and is related to clustering coefficients. Consequently, it has been applied in various fields ([4, 8, 13]). However, Ollivier–Ricci curvature cannot be defined on doubly-weighted graphs, and its computation can be expensive owing to the complexity involved. Conversely, the Forman–Ricci curvature relies on combinatorial principles and plays a vital role in identifying key vertices, akin to the Ollivier–Ricci curvature. Moreover, owing to its simpler definition, Forman curvature is generally computationally more efficient than the Ollivier–Ricci curvature. Thus, Forman–Ricci curvature has also been applied in various fields, including the identification of bottlenecks in networks ([6, 16]). Through comparative analysis on various networks, Samal et al. [14] demonstrated that both Ollivier–Ricci and Forman–Ricci curvatures can effectively identify important vertices. Furthermore, they demonstrated that Ollivier–Ricci curvature may be more suited for networks representing stochastic phenomena, whereas Forman–Ricci curvature might be more effective for networks reflecting actual distances. Crucially, Forman–Ricci curvature can be readily defined and computed for doubly-weighted graphs. Given our goal of developing a versatile analysis method applicable to doubly-weighted multiplex networks, we employed the Forman–Ricci curvature in this study. Henceforth, we will refer to it simply as Forman curvature.
The remainder of this paper is organized as follows: Section 2 defines the multiplex graph framework used, including relevant terminology. Section 3 introduces our definition of Forman curvature for multiplex graphs, establishes its theoretical properties, and illustrates its calculation with concrete examples. In Section 4, we introduce a new analysis method for multilayer networks and calculate some mathematical models. Finally, Section 5 summarizes our conclusions and discusses potential directions for future research.
2. Preliminaries
A graph is typically represented as a pair, $G=(V,E)$, where $V$ is a set of vertices and $E$ is a set of edges connecting pairs of vertices. In this basic representation, all edges are considered equivalent. However, in many cases, edges have distinct values or significance. For example, in a transportation network, an edge might represent an actual distance or transportation cost. To capture this, an edge-weight function $w:E\to\mathbb{R}$ is introduced. A graph equipped with such a function, denoted $V=(G,E,w)$, is called an edge-weighted graph. Alternatively, assigning values to vertices rather than edges can provide a more accurate representation in certain contexts. For example, in a transaction network, a vertex weight could represent the market capitalization or size of a company. Similarly, a vertex-weight function $m:V\to\mathbb{R}$ can be defined. A graph with this function, denoted $G=(V,E,m)$, is called a vertex-weighted graph. A graph incorporating both types of weights, denoted as $G=(V,E,w,m)$, is termed a doubly-weighted graph.
To integrate these graphs and represent them as multilayer graphs, it is necessary to add a new element, i.e., a set of “layers,” to this graph representation.
Definition 2.1.
A multiplex graph s defined by the structure $M=(V_{M},E_{M},V,\mathcal{L})$, where
  1. (1)
    $V=\left\{u_{1},u_{2},\cdots,u_{n}\right\}$ is the set of state vertices,
  2. (2)
    =$\mathcal{L}=\left\{1,\cdots,L\right\}$ is the set of layers,
  3. (3)
    $V_{M}=V\times\mathcal{L}$ is the set of vertices in each layer. Each element ($x^{i}$) in $V_{M}$ represents the vertex $x$ in layer $i$, and the subset of state vertices belonging to a specific layer $i$ is denoted by $V_{M}^{i}=V\times\left\{i\right\}$.
  4. (4)
    $E_{M}\subseteq V_{M}\times V_{M}$ is the set of edges, and $E_{M}$ is typically partitioned into a set of intra-layer edges, defined as $E_{A}:=\left\{(x^{i},y^{j})\in E_{M}\mid i=j\right\}$, and a set of inter-layer edges, defined as $E_{C}:=\left\{(x^{i},y^{j})\in E_{M}\mid x=y\right\}$. The set of edges in layer $i$ is denoted by $E_{M}^{i}$.
Specifically, when $L=1$, $M$ is referred to as a monolayer graph and corresponds to a standard graph.
Remark 2.2.
As noted by Kivelä et al. [10], the term multilayer graph encompasses various constructions for integrating several graph layers, and specific terminology often depends on the characteristics of the graphs to be integrated and the method of integration. This study focuses on multiplex graphs, characterized by having the same vertex set in each layer and inter-layer edges that connect only corresponding vertices across different layers.
Figure 1. Example of a multiplex graph $M=(V_{M},E_{M},V,\mathcal{L})$, where $V=\left\{0,1,2,3,4,5,6\right\}$, =$\mathcal{L}=\left\{1,2\right\}$, and $E_{M}$ comprises 11 edges.
Similar to monolayer graphs, weight functions can be defined for multiplex graphs. Specifically, $w:E_{M}\to\mathbb{R}$ is an edge-weight function and $m:V_{M}\to\mathbb{R}$ is a vertex-weight function. A multiplex graph, denoted by $M=(V_{M},E_{M},V,\mathcal{L},w,m)$, is called a doubly-weighted multiplex graph.
Henceforth, unless stated otherwise, the term “multiplex graph” in this paper will refer to a doubly-weighted multiplex graph.
We now adapt fundamental graph concepts to the multiplex setting.
Definition 2.3.
Let $M=(V_{M},E_{M},V,\mathcal{L},w,m)$ be a multiplex graph.
  1. (1)
    For any state vertex $x^{i}\in V_{M}$, the neighborhood of $x^{i}$ is defined by
    $\Gamma(x^{i})$ $=$ $\Gamma_{A}(x^{i})\cup\Gamma_{C}(x^{i})$
    ::=:= $\left\{y^{i}\in V_{M}\mid(x^{i},y^{i})\in E_{A}\right\}\cup\left\{x^{j}\in V_{M}\mid(x^{i},x^{j})\in E_{C}\right\}.$
    We also use the notation $\Gamma^{i,j}_{C}(x):=\Gamma_{C}(x^{i})\setminus\left\{x^{j}\right\}=\Gamma_{C}(x^{j})\setminus\left\{x^{i}\right\}$.
  2. (2)
    For any state vertex $x^{i}\in V_{M}$, the degree of $x^{i}$ is defined as
    $\deg(x^{i})=\sum_{y^{i}\in\Gamma(x^{i})}w((x^{i},y^{i})).$
When constructing a multiplex graph by combining multiple doubly-weighted graphs, the intra-layer weight of each edge of $E_{A}$ can be directly inherited from the original graph. However, it is important to define the weights of each inter-layer edge. Therefore, we introduce the concept of a compile graph. For this purpose, we first define the following quantity for any state vertex $x^{i}\in V_{M}$:
$W(x^{i}):=\begin{cases}\left(\sum_{y^{i}\in\Gamma_{A}(x^{i})}\cfrac{1}{\sqrt{w((x^{i},y^{i}))}}\right)^{-1},&\text{if}\ \Gamma_{A}(x^{i})\neq\emptyset,\\ 0,&\text{if}\ \Gamma_{A}(x^{i})=\emptyset.\end{cases}$
Definition 2.4.
Let $\left\{G_{l}=(V,E_{l},w_{l},m_{l})\right\}_{l=1}^{L}$ be a sequence of doubly-weighted graphs with the same number of vertices. A compile graph $CG$ is then defined as the multiplex graph
$V(CG)$ $=$ $V,\ \mathcal{L}=\left\{1,\cdots,L\right\},\ V_{CG}=\bigsqcup_{l}(V(G_{l})\times\left\{l\right\}),$
$E_{A}$ ::=:= $\left\{(x^{i},y^{j})\in V_{CG}\times V_{CG}\mid i=j,\ (x,y)\in E(G_{i})\right\},$
$E_{C}$ ::=:= $\left\{(x^{i},x^{j})\mid x\in V,\ i,j\in\mathcal{L},\ i\neq j\right\},$
where $m(x^{i})=m_{i}(x)$ for any vertex $x^{i}\in V_{CG}$, $w((x^{i},y^{i}))=w_{i}((x,y))$ for any intra-layer edge $(x^{i},y^{i})\in E_{A}$, and
$w((x^{i},x^{j}))=\min\left\{W^{2}_{i}(x),W^{2}_{j}(x)\right\}$
for any inter-layer edge $(x^{i},x^{j})\in E_{C}$.
Remark 2.5.
If all intra-layer edge weights equal 1, then $W(x^{i})=1/|\Gamma_{A}(x^{i})|$. In this case, the smaller the number of edges connected to a vertex $x^{i}$, the larger the value of $W(x^{i})$. To further investigate the behavior of $W(x^{i})$ in more detail when intra-layer edge weights are not equal to 1, we consider the function $f(x,y)=1/\sqrt{x}+1/\sqrt{y}$. If $x>0$, then
$\frac{\partial f}{\partial x}=-\frac{1}{x^{3/2}}<0.$
Thus, $f(x,y)$ is a monotonically decreasing function when $x>0$ and $y>0$. This is because the partial derivatives of $y$ are obtained similarly. As $W(x^{i})$ can be expressed as an inverse function of $f(x,y)$, $W(x^{i})$ monotonically increases with respect to each incident intra-layer edge weight. That is, $W(x^{i})<W(x^{j})$ suggests that the overall weight associated with the intra-layer edges connected to vertex $x^{j}$ is greater than that for those connected to vertex $x^{i}$.
3. Forman Curvature
This section defines Forman curvature, the specific type of discrete curvature employed in this study.
3.1. Forman curvature on graphs
This subsection introduces the definition of Forman curvature for standard (monolayer) graphs and outlines its key characteristics.
Definition 3.1 (Forman [7]).
Let $G=(V,E,w,m)$ be a doubly-weighted graph. The Forman curvature of any edge $e=(x,y)$ is defined as
$F((x,y))=2(m(x)+m(y))-m(x)\sum_{(x,z)\in E}\sqrt{\frac{w((x,y))}{w((x,z))}}-m(y)\sum_{(y,w)\in E}\sqrt{\frac{w((x,y))}{w((y,w))}}.$
Remark 3.2.
Key properties of Forman curvature include ([14, 15]):
  1. (1)
    Typically yields negative values for most edges.
  2. (2)
    Exhibits little correlation with clustering coefficients, but often shows a negative correlation with node degree and centrality measures.
  3. (3)
    Influences large-scale network connectivity. Removing nodes associated with low average Forman curvature can lead to faster network fragmentation than removing nodes based solely on high clustering coefficients. This suggests that nodes with a small Forman curvature are the most important for network integrity.
  4. (4)
    Unlike some traditional metrics for network structure analysis, such as centrality and clustering coefficient, which are primarily defined for unweighted or only edge-weighted graphs, Forman curvature can be defined for doubly-weighted graphs, making it suitable for the graphs considered in this study.
  5. (5)
    In unweighted graphs, Forman curvature is negatively correlated with both centrality and clustering coefficients (see Table 3 in [14]).
3.1.1. Example
To illustrate the relationship between Forman curvature and graph structure, we computed it for specific graphs. For simplicity, we consider unweighted graphs ($w=1$ and $m=1$). The Forman curvature is then computed as follows:
$F((x,y))=4-\deg(x)-\deg(y).$
Example 3.3.
Let $G=K_{n}$ be a complete graph on $n$ vertices.
For any vertex $x\in V$, $\deg(x)=n-1$. Thus, the Forman curvature of any edge $e=(x,y)$ is
$F((x,y))=4-2(n-1).$
Example 3.4.
Let $G=C_{n}$ be a cycle graph on $n$ vertices.
For any vertex $x\in V$, $\deg(x)=2$. Thus, the Forman curvature of any edge is zero.
Example 3.5.
Let $G=T_{r}$ be a $r$-regular tree graph.
For any vertex $x\in V$, we have $\deg(x)=r$. Thus, the Forman curvature of any edge $e=(x,y)$ is
$F((x,y))=4-2r.$
Figure 2. Example graphs (from left to right: complete, cycle, and tree)
As suggested by the definition of Forman curvature, the greater the number of edges connected to the edge, the more negative the value of the Forman curvature. Therefore, the smaller the value of the Forman curvature, the more important the edge is in the network.
3.2. Forman curvature on compile graphs
The definition of Forman curvature can be easily extended to multiplex graphs, provided vertex and edge weights are defined across all state vertices and edges.
Definition 3.6.
Let $M=(V_{M},E_{M},V,\mathcal{L},w,m)$ be a multiplex graph. For any edge $e=(x^{i},y^{j})\in E_{M}$, the Forman curvature is
$F(e)=2(m(x^{i})+m(y^{j}))-m(x^{i})\sum_{\bar{x}^{k}\in\Gamma(x^{i})}\sqrt{\frac{w(e)}{w(x^{i},\bar{x}^{k})}}-m(y^{j})\sum_{\bar{y}^{l}\in\Gamma(y^{j})}\sqrt{\frac{w(e)}{w(y^{j},\bar{y}^{l})}}.$
Specifically, for an intra-layer edge $e=(x^{i},y^{i})\in E_{A}$,
$F(e)=F^{i}((x,y))-m(x^{i})\sum_{x^{k}\in\Gamma_{C}(x^{i})}\sqrt{\frac{w((x^{i},y^{i}))}{w((x^{i},x^{k}))}}-m(y^{j})\sum_{y^{k}\in\Gamma_{C}(y^{i})}\sqrt{\frac{w((x^{i},y^{i}))}{w((y^{i},y^{k}))}},$
where $F^{i}$ is the Forman curvature on the $i$-th graph.
For the specific case of compile graphs, we introduce the following notations:
$\Gamma^{i,j}_{C,-}(x):=\left\{x^{k}\in\Gamma^{i,j}_{C}(x)\mid W(x^{k})\leq\min\left\{W(x^{i}),W(x^{j})\right\}\right\},$
$\Gamma^{i,j}_{C,+}(x):=\left\{x^{k}\in\Gamma^{i,j}_{C}(x)\mid W(x^{k})\geq\max\left\{W(x^{i}),W(x^{j})\right\}\right\}.$
The Forman curvature of any inter-layer edge can then be simplified as follows:
Definition 3.7.
Let $CG=(V_{CG},E_{CG},V,\mathcal{L},w,m)$ be a compile graph. The Forman curvature of any inter-layer edge $e=(x^{i},x^{j})\in E_{C}$ with $W(x^{i})\leq W(x^{j})$is expressed as
$F((x^{i},x^{j}))$ $=$ $m(x^{i})\left\{1-W(x^{i})\sum_{x^{k}\in\Gamma_{C}(x^{i})}\left(\frac{1}{W(x^{i})}\vee\frac{1}{W(x^{k})}\right)\right\}$
$+$ $m(x^{j})\left\{2-\frac{W(x^{i})}{W(x^{j})}-W(x^{i})\sum_{x^{l}\in\Gamma_{C}(x^{j})}\left(\frac{1}{W(x^{j})}\vee\frac{1}{W(x^{l})}\right)\right\}$
$=$ $-m(x^{i})|\Gamma_{C}^{i,j}(x)\setminus\Gamma_{C,-}^{i,j}(x)|+m(x^{j})\left(1-(|\Gamma_{C,+}^{i,j}(x)|+1)\cfrac{W(x^{i})}{W(x^{j})}\right)$
$-m(x^{i})\sum_{x^{k}\in\Gamma_{C,-}^{i,j}(x)}\cfrac{W(x^{i})}{W(x^{k})}-m(x^{j})\sum_{x^{l}\in\Gamma^{i,j}_{C}(x)\setminus\Gamma_{C,+}^{i,j}(x)}\cfrac{W(x^{i})}{W(x^{l})}$
where for real numbers $s$ and $t$, $s\wedge t:=\min\left\{s,t\right\}$ and $s\vee t:=\max\left\{s,t\right\}$.
We now establish some properties of the Forman curvature in multiplex graphs.
Theorem 3.8.
Let $CG=(V_{CG},E_{CG},V,\mathcal{L},w,m)$ be a compile graph. For any inter-layer edge $e=(x^{i},x^{j})\in E_{C}$ with $W(x^{i})\leq W(x^{j})$, we then have
(1) $F((x^{i},x^{j}))$$\geq$ $-(m(x^{i})+m(x^{j}))|\Gamma^{i,j}_{C,-}(x)|\max\left\{\cfrac{W(x^{i})}{W(x^{l})}\mid x^{l}\in\Gamma^{i,j}_{C,-}(x)\right\}$
$-m(x^{j})\left(\cfrac{W(x^{i})}{W(x^{j})}-1\right)$
(2) $\geq$ $-(m(x^{i})+m(x^{j}))|\Gamma^{i,j}_{C,-}(x)|\max\left\{\cfrac{W(x^{i})}{W(x^{l})}\mid x^{l}\in\Gamma^{i,j}_{C,-}(x)\right\}.$
Equality holds in the first inequality line if and only if $\Gamma^{i,j}_{C}(x)=\Gamma^{i,j}_{C,-}(x)$. Equality holds in the second inequality line if and only if $W(x^{i})=W(x^{j})$.
Proof.
Note that for any subset $A$ of $X$, if the number of elements in $A$ decreases by one, the number of elements in $X\setminus A$ increases by one.
If there exists a vertex $x^{k}\in V_{CG}$ such that $x^{k}\notin\Gamma_{C,-}^{i,j}(x)$, we consider another compile graph $\bar{CG}$ where only the value of $W(x^{k})$ is modified to ensure that $x^{k}$ is included in $\Gamma_{C,-}^{i,j}(x)$. Let the modified value of $W(x^{k})$ be $\bar{W}(x^{k})$. In this case, the difference in Forman curvature is calculated as follows:
$F^{CG}((x^{i},x^{j}))-F^{\bar{CG}}((x^{i},x^{j}))=m(x^{i})\left(\cfrac{W(x^{i})}{\bar{W}(x^{k})}-1\right)\geq 0,$
which implies $F^{CG}((x^{i},x^{j}))\geq F^{\bar{CG}}((x^{i},x^{j}))$. Thus, the larger the cardinality of $\Gamma_{C,-}^{i,j}(x)$, the smaller the Forman curvature.
Conversely, if there exists a vertex $x^{k}\in V_{CG}$ such that $x^{k}\notin\Gamma_{C,+}^{i,j}(x)$, we consider another compile graph $\hat{CG}$ where only the value of $W(x^{k})$ is modified to ensure that $x^{k}$ is included in $\Gamma_{C,+}^{i,j}(x)$. Let the modified value of $W(x^{k})$ be $\hat{W}(x^{k})$. In this case, the difference in Forman curvature is calculated as follows:
$F^{CG}((x^{i},x^{j}))-F^{\hat{CG}}((x^{i},x^{j}))=m(x^{j})\left(\cfrac{W(x^{i})}{W(x^{j})}-\cfrac{W(x^{i})}{W(x^{k})}\right)\leq 0,$
which implies $F^{CG}((x^{i},x^{j}))\leq F^{\hat{CG}}((x^{i},x^{j}))$. Thus, the larger the cardinality of $\Gamma_{C,+}^{i,j}(x)$, the larger the Forman curvature.
Based on these results, the minimum value of the Forman curvature occurs when the cardinality of $\Gamma_{C,-}^{i,j}(x)$ coincides with that of $\Gamma_{C}^{i,j}(x)$. The minimum value is calculated as follows:
$F((x^{i},x^{j}))$$\geq$ $-(m(x^{i})+m(x^{j}))|\Gamma^{i,j}_{C,-}(x)|\max\left\{\cfrac{W(x^{i})}{W(x^{l})}\mid x^{l}\in\Gamma^{i,j}_{C,-}(x)\right\}$
$-m(x^{j})\left(\cfrac{W(x^{i})}{W(x^{j})}-1\right).$
Next, we examine the maximum value of $W(x^{i})/W(x^{j})$. Given the condition that $W(x^{i})\leq W(x^{j})$, the maximum value of $W(x^{i})/W(x^{j})$ is 1 when $W(x^{i})=W(x^{j})$. Thus, the lower bound of the Forman curvature is proved. ∎
Theorem 3.9.
Let $CG=(V_{CG},E_{CG},V,\mathcal{L},w,m)$ be a compile graph. Then for any inter-layer edge $e=(x^{i},x^{j})\in E_{C}$ with $W(x^{i})\leq W(x^{j})$, the Forman curvature of $e$ is represented as follows.
$F((x^{i},x^{j}))\leq-|\Gamma^{i,j}_{C}(x)|m(x^{i})+m(x^{j})\left(1-(|\Gamma_{C}^{i,j}(x)|+1)\cfrac{W(x^{i})}{W(x^{j})}\right).$
The condition for equality is $\Gamma^{i,j}_{C}(x)=\Gamma^{i,j}_{C,+}(x)$.
Proof.
In the proof of Theorem 3.8, we established the following relationships:
  1. (1)
    The smaller the cardinality of $\Gamma_{C,-}^{i,j}(x)$, the larger the Forman curvature.
  2. (2)
    The larger the cardinality of $\Gamma_{C,+}^{i,j}(x)$, the larger the Forman curvature.
Thus, the Forman curvature attains its maximum value when $\Gamma^{i,j}_{C}(x)=\Gamma^{i,j}_{C,+}(x)$. The maximum value is computed as follows:
$F((x^{i},x^{j}))\leq-|\Gamma^{i,j}_{C}(x)|m(x^{i})+m(x^{j})\left(1-(|\Gamma_{C}^{i,j}(x)|+1)\cfrac{W(x^{i})}{W(x^{j})}\right).$
This completes the proof. ∎
3.2.1. Example
To illustrate further, we compile graphs with two or three layers and calculate the Forman curvature of the representative inter-layer and intra-layer edges.
Example 3.10.
Let $CG=(V_{CG},E_{CG},V,\mathcal{L},w,m)$ be a compile graph with $L=2$.
The Forman curvature of any inter-layer edge $e=(x^{1},x^{2})\in E_{C}$ with $W(x^{1})\leq W(x^{2})$ is calculated as
$F((x^{1},x^{2}))=m(x^{2})\left(1-\cfrac{W(x^{1})}{W(x^{2})}\right)\geq 0.$
Conversely, the Forman curvature of any intra-layer edge $e=(x^{1},y^{1})\in E_{A}$ is given by
$F(e)=F^{1}((x,y))-m(x^{1})\frac{\sqrt{w(e)}}{\min\left\{W(x^{1}),W(x^{2})\right\})}-m(y^{1})\cfrac{\sqrt{w(e)}}{\min\left\{W(y^{1}),W(y^{2})\right\})}.$
Example 3.11.
Let $CG=(V_{CG},E_{CG},V,\mathcal{L},w,m)$ be a compile graph with $L=3$.
The Forman curvature of any inter-layer edge $e=(x^{1},x^{2})\in E_{C}$ with $W(x^{1})\leq W(x^{2})$ is then expressed as
$F(e)=-m(x^{1})\left(1\vee\cfrac{W(x^{1})}{W(x^{3})}\right)+m(x^{2})\left(1-\cfrac{W(x^{1})}{W(x^{2})}-\cfrac{W(x^{1})}{W(x^{2})}\vee\cfrac{W(x^{1})}{W(x^{3})}\right)$
$=$ $\begin{cases}-m(x^{1})\cfrac{W(x^{1})}{W(x^{3})}+m(x^{2})(1-\cfrac{W(x^{1})}{W(x^{2})}-\cfrac{W(x^{1})}{W(x^{3})}),&\text{if}\ W(x^{3})\leq W(x^{1})\leq W(x^{2}),\\ -m(x^{1})+m(x^{2})(1-\cfrac{W(x^{1})}{W(x^{2})}-\cfrac{W(x^{1})}{W(x^{3})}),&\text{if}\ W(x^{1})<W(x^{3})<W(x^{2}),\\ -m(x^{1})+m(x^{2})(1-2\cfrac{W(x^{1})}{W(x^{2})}),&\text{if}\ W(x^{1})\leq W(x^{2})\leq W(x^{3}).\end{cases}$
Conversely, the Forman curvature of any intra-layer edge $e=(x^{1},y^{1})\in E_{A}$ is given by
$F(e)$ $=$ $F^{1}((x,y))-m(x^{1})\frac{\sqrt{w(e)}}{\min\left\{W(x^{1}),W(x^{2})\right\})}-m(x^{1})\frac{\sqrt{w(e)}}{\min\left\{W(x^{1}),W(x^{3})\right\})}$
$-m(y^{1})\cfrac{\sqrt{w(e)}}{\min\left\{W(y^{1}),W(y^{2})\right\})}-m(y^{1})\cfrac{\sqrt{w(e)}}{\min\left\{W(y^{1}),W(y^{3})\right\})}.$
These examples demonstrate that the Forman curvature of an inter-layer edge $e=(x^{i},x^{j})$ in a compile graph depends on whether each $W(x^{k})\in\Gamma_{C}^{i},j(x)$ belongs to $\Gamma_{C,+}^{i},j(x)$ or $\Gamma_{C,-}^{i},j(x)$. Additionally, the Forman curvature of an intra-layer edge is highly influenced by the Forman curvature of the $i$-graph and can serve as a measure of edge importance within the $i$-graph, as well as the Forman curvature of the graph.
4. Graph Algorithm
If the primary goal when analyzing a multilayer network is to identify vertices that maintain consistent importance across all layers, calculating importance within each layer separately may suffice. This approach can reduce the need for a dedicated multilayer framework. Conversely, identifying vertices whose importance varies between layers and minimizing these differences requires considering the multilayer network as a whole. For example, identifying a location within a transportation network crucial for one public transit system but marginal for another can highlight specific points warranting focus for future network development.
4.1. Structure
Before outlining the algorithm steps, we define a new indicator based on inter-layer Forman curvature.
Definition 4.1.
Let $CG=(V_{CG},E_{CG},V,\mathcal{L},w,m)$ be a compile graph. For a vertex $x\in V$, a comprehensive evaluation of $x$ is defined as
$CE(x)=\sum_{i,j\in\mathcal{L}:i<j}F((x^{i},x^{j})).$
Remark 4.2.
For a vertex $x\in V$, consider the layers ordered such that $W(x_{1})\leq W(x_{2})\leq\cdots\leq W(x_{L})$. Subsequently, applying Theorem 3.8, a ower bound for the comprehensive evaluation of $x$ is represented as follows:
$CE(x)$ $=$ $\sum_{i,j\in\mathcal{L}:i<j}F((x^{i},x^{j}))$
$\geq$ $-\sum_{i,j\in\mathcal{L}:i<j}(m(x^{i})+m(x^{j}))|\Gamma^{i,j}_{C,-}(x)|\max\left\{\cfrac{W(x^{i})}{W(x^{l})}\mid x^{l}\in\Gamma^{i,j}_{C,-}(x)\right\}$
$=$ $-\sum_{i,j\in\mathcal{L}:i<j}(m(x^{i})+m(x^{j}))|\Gamma^{i,j}_{C,-}(x)|\cfrac{W(x^{i})}{W(x^{1})}.$
Alternatively, for the case of $W(x^{1})=\cdots=W(x^{L})$, the comprehensive evaluation of $x$ is denoted by $CE^{\text{uni}}(x)$ and calculated as
$CE^{\text{uni}}(x)$ $=$ $\sum_{i,j\in\mathcal{L}}F((x^{i},x^{j}))$
$=$ $-\sum_{i,j\in\mathcal{L}}(m(x^{i})+m(x^{j})).$
Therefore, for a given vertex $x\in V$, a larger difference between $CE(x)$ and $CE^{\text{uni}}(x)$ suggests greater dispersion among the $W(x^{i})$ values for the vertex. Note that while variance and standard deviation can be used as indicators of variation, they might not be suitable for comparing the $W(x^{i})$ dispersion across different vertices, as the representative value is different for each vertex.
Algorithm 1.
Input a compile graph $CG=(V_{CG},E_{CG},V,\mathcal{L},w,m)$.
  1. Step 1.
    Calculate the difference between $CE(x)=\sum_{i<j}F((x^{i},x^{j}))$ and $CE^{\text{uni}}(x)$ for each vertex.
  2. Step 2.
    Compute $\sum_{y}F((x_{0}^{i},y^{i}))$ for each layer $i$ of the vertex $x_{0}$ identified as having a large ”Difference value” in Step 1.
  3. Step 3.
    Compute $F((x_{0}^{i_{0}},y^{i_{0}}))$ for each intra-layer edge $e=(x_{0}^{i_{0}},y^{i_{0}})$ for a layer $i_{0}$ with identified in Step 2 as having a large value and extract an edge $e_{0}=(x_{0}^{i_{0}},y^{i_{0}})$ that exhibits a large curvature value.
First, Step 1 calculates the difference between $CE(x)$ and $CE^{\text{uni}}(x)$ for each vertex $x$. This allows for a comparison of the degree of variability among the $W(x^{i})$ values for that vertex (as discussed in Remark 4.2). Next, Step 2 examines a vertex $x_{0}$, identified as having a large difference value in Step 1. By comparing the sum of incident intra-layer edge Forman curvatures for each layer $i$, this step helps distinguish layers where $x_{0}$ is important from those where it is not important (see Remark 3.2). Finally, Step 3 focuses on a specific layer $i_{0}$ where $x_{0}$ is not important by comparing the Forman curvature of each edge.
4.2. Example
This section describes the application of this graph algorithm to compile graphs using the classical Erdös–Rényi random graphs $G(n,p)$. This graph has $n$ vertices, and each possible edge between pairs of vertices appears independently with probability $p$. Additionally, a random number between 0.01 and 1 is assigned as the weight for each vertex. Similarly, a random number between 1 and 10 is assigned as the weight of each edge. These random graphs are then combined to construct a compile graph (Figure 3).
Figure 3.
Left: $CG^{258}$ (Compile graph of $\left\{G(25,0.2),G(25,0.5),G(25,0.8)\right\}$).
Right: $CG^{888}$ (Compile graph of $\left\{G(25,0.8),G(25,0.8),G(25,0.8)\right\}$).
Example 4.3.
$CG=CG^{258}$ is the compile graph constructed from
$\left\{G(25,0.2),G(25,0.5),G(25,0.8)\right\}$.
The results from applying this graphing algorithm show that Vertex 10 has the largest variation in its $W(x^{i})$ values across the layers (Figure 4). Moreover, the sum of the intra-layer edge Forman curvatures for Vertex 10, when calculated for Layers 1, 2, and 3, is largest in Layer 1. This leads to the conclusion that Vertex 10 is least important in Layer 1 (Figure 5). Finally, within Layer 1, the Forman curvature for each edge between Vertices 10 and 20 has the largest curvature value (Figure 6).
Figure 4. Results of running Step 1 for $CG^{258}$. The ”Difference” value represents the difference between $CE(x)$ and $CE^{\text{uni}}(x)$.
Figure 5. Results of running Step 2 for $CG^{258}$.
Figure 6. Results of running Step 3 for Vertex 10 and Layer 1 in $CG^{258}$.
Example 4.4.
$CG=CG^{888}$ is the compile graph constructed from
$\left\{G(25,0.8),G(25,0.8),G(25,0.8)\right\}$.
The results from this graphing algorithm indicate that Vertex 8 exhibits the largest in its $W(x^{i})$ value across the layers (Figure 7). Additionally, the sum of the Forman curvatures of the intra-layer edges for Vertex 8 across Layers 1, 2, and 3 reveals that the largest value occurs in Layer 3. This leads to the conclusion that Vertex 10 is least important in Layer 3 (Figure 8). Finally, within Layer 3, calculation of the Forman curvature for each edge connected to Vertex 8 shows that the edge between Vertices 8 and 22 has the largest curvature value (Figure 9).
Figure 7. Results of running Step 1 for $CG^{888}$.
Figure 8. Results of running Step 2 for $CG^{888}$.
Figure 9. Results of running Step 3 for Vertex 8 and Layer 3 in $CG^{888}$.
Comparing the two examples shows that when the graph structure of each layer is similar, the overall ”Difference value” tends to be smaller for all vertices $x$, and the variance of these difference values across vertices is also smaller. Conversely, when the graph structures of the layers differ significantly, the ”Difference values” are generally larger, and their variance across vertices is also larger. Therefore, comparing these difference values is useful not only for clarifying the degree of variation in $W(x^{i})$ for each vertex but also for understanding the graph structure of the entire multilayer network.
5. Conclusion
This paper introduced the Forman curvature for multilayer networks and demonstrated its utility. Specifically, we showed that computing the curvature of intra-layer edges allows for assessing vertex/edge importance within individual layers, while computing that of inter-layer edges provides insights into the relationships between layers and aspects of the overall multilayer network structure. The analysis of intra-layer curvature confirms that Forman curvature can play a similar role in assessing local importance within layers as in single-layer graphs. Concurrently, the analysis based on inter-layer curvature represents a novel application, offering a new method to use the Forman curvature. Although various centrality measures exist for multilayer networks and serve as important tools for assessing node importance, many are not applicable to the doubly-weighted multiplex graphs considered in this study. Therefore, the Forman-curvature-based indices and the analysis framework introduced herein hold potential for application across a wide range of fields dealing with complex, weighted multilayer systems. In future research, we plan to investigate the properties of inter-layer Forman curvature in greater detail and quantify the degree of variation of $W(x^{i})$ with greater precision.
Acknowledgment
We would like to thank Editage (www.editage.jp) for English language editing.
Funding
This work was supported in part by JSPS KAKENHI [grant number 25K17253].
Conflict of interest
None.